# trans-subtitle [https://orcid.org/0000-0003-0875-8345] [https://orcid.org/0000-0002-6003-3168] [https://orcid.org/0009-0000-1789-4117] [https://orcid.org/0009-0000-8135-7260] [https://orcid.org/0000-0003-4632-5743] [https://orcid.org/0009-0000-5184-2695] [https://orcid.org/0000-0003-1324-1412]

### v1

@@
Abstract
[ABSTRACT] Deploying nitrogen vacancy (NV) centers in diamond as nanoscale quantum sensors for condensed matter and materials physics requires placing the NV centers close to the sensing target. One solution is to fabricate diamond nanostructures and integrate them with materials and devices. However, diamond etching and ion milling can introduce subsurface damage and surface defects that degrade the charge stability and spin coherence of NV centers near the surface. Here we report a procedure for fabricating low-damage nanostructured diamond membranes, and we show that this fabrication scheme preserves the optical and spin properties of state-of-the-art shallow NV center quantum sensors, within nanometers of the diamond surface, while providing significant photonic enhancement. Furthermore, we demonstrate a pick-and-place transfer method, which enables integration with diverse sensing targets.
©This is an open access article under the terms of the Creative Commons Attribution-NonCommercial License, which permits use, distribution and reproduction in any medium, provided the original work is properly cited and is not used for commercial purposes.
© 2024 The Author(s) Advanced Functional Materials published by Wiley Periodicals LLC on behalf of Advanced Functional Materials.
keywords:
quantum sensing | diamond | nanofabrication | nitrogen vacancy | photonics | membrane | nanobeam
journal: Advanced Functional Materialsvolume: 0footnotetext: *These authors contributed equally to this work.
@@ 1 Introduction
There have been several advances in recent years in diamond nanofabrication, including reactive ion etching to define membranes, 19, pillars20, 21, 14 and nanobeams, 22, 23, 24 and smart-cut techniques to fabricate diamond-on-insulator devices. 25, 26, 27 An alternative method for realizing free-standing nanostructures is a modified version of the single crystal reactive etching and metallization (SCREAM) process, which was first developed for silicon 28 and recently adapted to single crystal, monolithic diamond. 29, 30 In this process, anisotropic, top-down etching is followed by the creation of a sidewall hardmask and subsequent quasi-isotropic etching to undercut the desired structures. In the quasi-isotropic etching step, setting the RF bias to zero induces a highly diffuse, isotropic, plasma that etches different crystallographic directions of diamond at different rates. 31 This strategy has been used to fabricate microdisk resonators 32 and diamond nanobeams. 30, 33, 34, 35, 36
Here we apply the quasi-isotropic etch as part of a process to fabricate a membrane consisting of a dense array of nanobeams, with sacrificial tethers to the host substrate. To achieve a highly uniform undercut across microns of dense nanostructures, we design the device to reduce variations in the quasi-isotropic etch rate. We perform the etch at lower temperature and inductively coupled plasma reactive ion etching (ICP-RIE) power than in reference 32, reducing the etch rate and enabling control over the undercut geometry as well as the beam and tether thicknesses. In particular, the sacrificial tether geometry can be tuned to achieve robust suspension while still allowing mechanical separation. This enables device integration through pick-and-place transfer via micro-manipulator probes.
By simultaneously ensuring that the surfaces have low contamination, low subsurface damage, and high quality oxygen termination, we fabricate nanobeams that host NV centers within nanometers of the surface that have state-of-the-art nanoscale sensing performance. We show that the spin coherence time, charge state stability, and charge dynamics of shallow NV centers in these nanobeams are in line with "typical best" bulk substrate processing. Furthermore, the nanobeam form factor enables collection efficiency enhancement by up to a factor of seven when integrated with a sapphire target, in agreement with photonic simulations.
@@ 2.1 Diamond Nanobeam Membrane Fabrication
The fabrication process follows the approach of Mitchell et al. 32, with several adaptations to create high-density nanobeam membranes compatible with pick-and-place device integration for quantum sensing applications, as described below.
@@
We prepare the surface of a (100) oriented electronic grade diamond (Element Six), according to the process outlined in 18 and implant it with 15N (3 keV, 1e9 cm-2) followed by an 800C activation anneal in vacuum to create shallow NV centers. We then deposit 150 nm of silicon nitride (SiNx) at 250C using plasma-enhanced chemical vapor deposition (PE-CVD, Oxford PlasmaPro 100), and pattern an etch mask (Zeon ZEP 520a) with electron beam lithography (Elionix ELS-F125), using a layer of spin-on conductive polymer (DisChem, DisCharge H2O X4) to reduce charging effects. The SiNx hard mask is patterned with a highly anisotropic etch (Figure 1a) using inductively-coupled plasma reactive ion etching (ICP RIE, PlasmaTherm Takachi SLR, 14 sccm CHF3 and 10 sccm SF6 at 5 mTorr, 200 W RF and 60 W ICP power). The high selectivity of the nitride in oxygen plasmas enables very smooth fabricated structures (Figure 1b).
@@
We first fabricate the membranes with variable nanobeam widths (120 - 720 nm) to study how nanobeam width and thickness influence the fabrication procedure. Based on finite-difference time-domain modeling (FDTD) (Figure S.7 and Section 2.4), we target $\approx$700 nm width and $\approx$350 nm thickness, which sets the dimensions for all the features in a given fabrication run. Due to the upward etch rate of the quasi-isotropic etch, we etch to a deeper depth than the desired thickness during the anisotropic diamond etch, balancing these two etch rates to achieve the final desired thickness. Etching in pure oxygen plasma (10 mTorr, 110 W ICP-RIE, 1000 W RF Bias) for 165 s removes 820 nm of diamond with near vertical side walls, no edge collapse, and minimal micromasking (Figure S.1). Following the anisotropic etch, we deposit another 150 nm of SiNx conformally over the etched, SiNx-capped diamond mesa. The residual SiNx hard mask is thinned during the anisotropic etch step but is still continuous. The hard mask on the bottom diamond surface between the mesas is selectively removed using another directional SiNx etch process, leaving a SiNx hard mask layer on the sidewalls and top surface. Crucially, the resulting encapsulation of the corners of the device is sufficient to protect the top edge of the membrane during the quasi-isotropic etch.
@@
The critical step is the long, quasi-isotropic etch performed in the same ICP-RIE with the platen heated to 140 °C. After an initial RF strike, the etch proceeds for 20 hours with 1000 W ICP and 0 W RF at 20 mTorr, resulting in quasi-isotropic etching along multiple directions. The lateral etch rate is strongly related to the distance between structures, as the quasi-isotropic etch rate is highly aspect-ratio dependent, necessitating long etches to achieve full undercut. For a 1 µm spacing, we find that the quasi-isotropic etch etches $\approx$600 nm laterally in 20 hours, guaranteeing full undercut for the targeted nanobeam width range. This aspect-ratio dependent rate is much smaller than in previous reports for structures such as microdisks, e.g. 2000 nm in 10 hrs ,29 and of sparse density rectangular photonic crystal cavities, 36 which do not suffer from an aspect ratio throttled etch rate.
@@
Another key aspect enabling the pick-and-place operation is careful design of the outer frame (“trellis”) and attachment tethers. The double-layer trellis remains rigid, enabling transferable nanobeam frames with individual nanobeams as narrow as 100 nm and tens of micrometers long. The standardized trellis results in a high yield process, enabling consistent undercutting of thousands of membranes across a 4.5 x 4.5 mm2 sample(Figure 1e). The 1 µm gap between the frame and host crystal ensures that the etch rate of the nanobeams is uniform even at the edges of the pattern, and the trellis buffers against aspect-ratio dependent edge effects, particularly at the corners (Figure 1c) where there is significantly faster upwards etching during the quasi-isotropic step. The tethers are very narrow (~100 nm) and thin to ensure mechanical flexibility37, and we pattern a large number to build a mechanically reliable anchor. These tethers can be broken individually by micromanipulators while allowing the membranes to withstand harsh processing. Without this design, frames would collapse or delaminate, precluding scalable, high-yield fabrication.
@@
We use X-ray photoelectron spectroscopy to verify surface cleanliness prior to performing a thermal oxygen anneal at 450C according to the procedure described in 18, which ensures high-quality oxygen termination. We image the resulting NV centers in a confocal microscope (NA = 0.9) under 520 nm illumination. The confocal microscope image overlaid with a scanning electron microscopy (SEM) image of the devices shows bright, single NV center emission in the nanobeams (Figure 1f). Single NV centers are brighter in the nanobeams compared to the host diamond, while their density is similar.
Figure 2: (a) Hahn echo coherence time T2 as a function of NV center depth in nanobeams and the host diamond, plotted together with data from Ref 18 for shallow NV centers after the oxygen surface termination procedure including a 1200C and only a 800C anneal (as in the current work). Yellow dashed line shows a $T_{2}=kd^{n}+T_{2,0}$ fit to the 800C annealed data from Ref. 18 ($k=0.047\pm 0.04,n=2.35\pm 0.25,T_{2,0}=3.35\pm 3.87$). Grey shaded area marks the error margin for $n$ with fixed $k=0.047$ and $T_{2,0}=3.35$ showing that most of the data that we measure in the nanobeams falls within this scaling range. (b) T2 measured using an XY-N pulse sequence as a function of the number of pulses, N, (N = 1 is a Hahn echo experiment) for two NV centers in the nanobeams. Dashed line shows $\propto{N^{2/3}}$ dependence. (c) Spectral noise density extracted from a XY-8 decoherence decay using Equation S2 for the same NV centers as in (b). The expected peak position associated with the 13C Larmor frequency is marked with a vertical dashed line. (d) $T_{1,SQ}$ and $T_{1,DQ}$ as a function of NV center depth in nanobeams (purple) and unfabricated diamond (blue).
@@ 2.2 NV Center Spin Properties
We examine the spin properties of NV centers implanted within nanometers of the diamond surface in nanobeams to evaluate their utility as nanoscale quantum sensors, and we compare them to those in the unfabricated regions of the diamond. We measure the Hahn echo coherence time ($T_{2}$) as a function of depth from the surface for several NV centers in the nanobeams and in the unfabricated diamond (Figure 2a). The depth is measured using proton nuclear magnetic resonance, as has been described previously. 38 The $T_{2}$ decreases as NV centers are brought closer to the diamond surface, as has been reported in the literature. 18, 39 The observed depth dependence of $T_{2}$ for NV centers in nanobeams is within the distribution of NV centers in the unfabricated regions, and is also consistent with state-of-the-art surface termination as described in the literature. 18 We note that the coherence times are slightly shorter than the best reported NV centers (Figure 2a, orange points), which resulted from a 1200C anneal that removes multivacancy clusters that form because of implantation damage, pointing to a potential direction for future improvement in the NV center performance in nanobeams.
@@
To probe higher frequency noise, we measure the spin coherence with dynamical decoupling by XY-N pulse sequences. The coherence time $T_{2}$ increases with the number of pulses in the sequence, $N$, for two randomly selected NV centers (Figure 2b). We do not observe any saturation in $T_{2}$ up to 96 pulses, however the scaling deviates from the $N^{2/3}$} scaling expected for a slowly fluctuating noise bath,40 which can be attributed to the presence of broadband noise induced by defects at the diamond surface. Spectral noise decomposition using an XY-8 pulse sequence41 for the same two NV centers reveals a noise spectrum $S(\omega)$ consistent with a broadband noise bath and a pronounced peak corresponding to the 13C nuclear spin Larmor precession at 360 G (Figure 2c). These observations are consistent with previous characterization of the surface noise bath in state-of-the-art shallow NV centers. 18
@@
Finally, we measure single-quantum spin lifetime ($T_{1,SQ}$) and double-quantum spin lifetime ($T_{1,DQ}$) to probe magnetic and electric noise near the NV center spin resonance frequency ($\approx$ 2.8 GHz)39, correspondingly, as a function of distance from the surface (Figure 2d). The spin lifetimes display a large distribution from 100 µs to several ms, with no clear depth dependence, and the distributions between NV centers in nanobeams and in the unfabricated region are similar. The $T_{1,DQ}$ values between 100 µs and 1 ms are similar to the results presented in 18, which confirms that the fabrication process does not contribute additional sources of electric noise from the diamond surface. At the same time, we observe several shallow NV centers with $T_{1,SQ}$ in the range of 200 µs to 1 ms compared to $T_{1,SQ}$ > 1 ms as reported in 18, which could point to slightly increased high-frequency magnetic noise at the surface.

### v2

# High-quality nanostructured diamond membranes for nanoscale quantum sensing

@@
High-quality nanostructured diamond membranes for nanoscale quantum sensing
Alexander C Pakpour-Tabrizi These authors contributed equally to this work. Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Artur Lozovoi These authors contributed equally to this work. Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Sean Karg These authors contributed equally to this work. Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Tecla Bottinelli Montandon University of Basel, Department of Physics, Basel, Switzerland    Melody Leung Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Kai-Hung Cheng Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA    Nathalie P. de Leon npdeleon@princeton.edu Princeton University, Department of Electrical and Computer Engineering, Princeton, NJ, USA
(November 16, 2025)
Deploying nitrogen vacancy (NV) centers in diamond as nanoscale quantum sensors for condensed matter and materials physics requires placing the NV centers close to the sensing target. One solution is to fabricate diamond nanostructures and integrate them with materials and devices. However, diamond etching and ion milling can introduce subsurface damage and surface defects that degrade the charge stability and spin coherence of NV centers near the surface. Here we report a procedure for fabricating low-damage nanostructured diamond membranes, and we show that this fabrication scheme preserves the optical and spin properties of state-of-the-art shallow NV center quantum sensors, within nanometers of the diamond surface, while providing significant photonic enhancement. Furthermore, we demonstrate a pick-and-place transfer method, which enables integration with diverse sensing targets.
@@ 1 Introduction
There have been several advances in recent years in diamond nanofabrication, including reactive ion etching to define membranes, [corazza2025homogeneous], pillars[Wrachtrup_pillar, Maletinsky_AFMCryo, scheidegger2022scanning] and nanobeams, [Loncar_beams_fab, Yacoby_tribeams_chempot, Du_tribeams_Fe] and smart-cut techniques to fabricate diamond-on-insulator devices. [Membrane_2_High, guo_direct-bonded_2024, ding_high-q_2024] An alternative method for realizing free-standing nanostructures is a modified version of the single crystal reactive etching and metallization (SCREAM) process, which was first developed for silicon [SCREAM] and recently adapted to single crystal, monolithic diamond. [Khanaliloo_highQ, Barclay_waveguide] In this process, anisotropic, top-down etching is followed by the creation of a sidewall hardmask and subsequent quasi-isotropic etching to undercut the desired structures. In the quasi-isotropic etching step, setting the RF bias to zero induces a highly diffuse, isotropic, plasma that etches different crystallographic directions of diamond at different rates. [Yacoby_isodeepdive] This strategy has been used to fabricate microdisk resonators [Barclay_300000] and diamond nanobeams. [Barclay_waveguide, TeonoFibreCoupled, hanson_waveguide, brevoord_large-range_2025, Englund_phc]
Here we apply the quasi-isotropic etch as part of a process to fabricate a membrane consisting of a dense array of nanobeams, with sacrificial tethers to the host substrate. To achieve a highly uniform undercut across microns of dense nanostructures, we design the device to reduce variations in the quasi-isotropic etch rate. We perform the etch at lower temperature and inductively coupled plasma reactive ion etching (ICP-RIE) power than in reference [Barclay_300000], reducing the etch rate and enabling control over the undercut geometry as well as the beam and tether thicknesses. In particular, the sacrificial tether geometry can be tuned to achieve robust suspension while still allowing mechanical separation. This enables device integration through pick-and-place transfer via micro-manipulator probes. By simultaneously ensuring that the surfaces have low contamination, low subsurface damage, and high quality oxygen termination, we fabricate nanobeams that host NV centers within nanometers of the surface that have state-of-the-art nanoscale sensing performance. We show that the spin coherence time, charge state stability, and charge dynamics of shallow NV centers in these nanobeams are in line with ”typical best” bulk substrate processing. Furthermore, the nanobeam form factor enables collection efficiency enhancement by up to a factor of seven when integrated with a sapphire target, in agreement with photonic simulations.
@@ 2.1 Diamond Nanobeam Membrane Fabrication
The fabrication process follows the approach of Mitchell et al. [Barclay_300000], with several adaptations to create high-density nanobeam membranes compatible with pick-and-place device integration for quantum sensing applications, as described below.
@@
We prepare the surface of a (100) oriented electronic grade diamond (Element Six), according to the process outlined in [Sangtawesin_SurfNoise] and implant it with 15N (3 keV, 1e9 cm-2) followed by an 800C activation anneal in vacuum to create shallow NV centers. We then deposit 150 nm of silicon nitride (SiNx) at 250C using plasma-enhanced chemical vapor deposition (PE-CVD, Oxford PlasmaPro 100), and pattern an etch mask (Zeon ZEP 520a) with electron beam lithography (Elionix ELS-F125), using a layer of spin-on conductive polymer (DisChem, DisCharge H2O X4) to reduce charging effects. The SiNx hard mask is patterned with a highly anisotropic etch (Figure 1a) using inductively-coupled plasma reactive ion etching (ICP RIE, PlasmaTherm Takachi SLR, 14 sccm CHF3 and 10 sccm SF6 at 5 mTorr, 200 W RF and 60 W ICP power). The high selectivity of the nitride in oxygen plasmas enables very smooth fabricated structures (Figure 1b).
@@
We first fabricate the membranes with variable nanobeam widths (120 - 720 nm) to study how nanobeam width and thickness influence the fabrication procedure. Based on finite-difference time-domain modeling (FDTD) (Figure S.7 and Section S.4), we target $\approx$700 nm width and $\approx$350 nm thickness, which sets the dimensions for all the features in a given fabrication run. Due to the upward etch rate of the quasi-isotropic etch, we etch to a deeper depth than the desired thickness during the anisotropic diamond etch, balancing these two etch rates to achieve the final desired thickness. Etching in pure oxygen plasma (10 mTorr, 110 W ICP-RIE, 1000 W RF Bias) for 165 s removes 820 nm of diamond with near vertical side walls, no edge collapse, and minimal micromasking (Figure S.1). Following the anisotropic etch, we deposit another 150 nm of SiNx conformally over the etched, SiNx-capped diamond mesa. The residual SiNx hard mask is thinned during the anisotropic etch step but is still continuous. The hard mask on the bottom diamond surface between the mesas is selectively removed using another directional SiNx etch process, leaving a SiNx hard mask layer on the sidewalls and top surface. Crucially, the resulting encapsulation of the corners of the device is sufficient to protect the top edge of the membrane during the quasi-isotropic etch.
@@
The critical step is the long, quasi-isotropic etch performed in the same ICP-RIE with the platen heated to 140 °C. After an initial RF strike, the etch proceeds for 20 hours with 1000 W ICP and 0 W RF at 20 mTorr, resulting in quasi-isotropic etching along multiple directions. The lateral etch rate is strongly related to the distance between structures, as the quasi-isotropic etch rate is highly aspect-ratio dependent, necessitating long etches to achieve full undercut. For a 1 µm spacing, we find that the quasi-isotropic etch etches $\approx$600 nm laterally in 20 hours, guaranteeing full undercut for the targeted nanobeam width range. This aspect-ratio dependent rate is much smaller than in previous reports for structures such as microdisks, e.g. 2000 nm in 10 hrs ,[Khanaliloo_highQ] and of sparse density rectangular photonic crystal cavities, [Englund_phc] which do not suffer from an aspect ratio throttled etch rate.
@@
Another key aspect enabling the pick-and-place operation is careful design of the outer frame (“trellis”) and attachment tethers. The double-layer trellis remains rigid, enabling transferable nanobeam frames with individual nanobeams as narrow as 100 nm and tens of micrometers long. The standardized trellis results in a high yield process, enabling consistent undercutting of thousands of membranes across a 4.5 x 4.5 mm2 sample(Figure 1e). The 1 µm gap between the frame and host crystal ensures that the etch rate of the nanobeams is uniform even at the edges of the pattern, and the trellis buffers against aspect-ratio dependent edge effects, particularly at the corners (Figure 1c) where there is significantly faster upwards etching during the quasi-isotropic step. The tethers are very narrow (~100 nm) and thin to ensure mechanical flexibility[Suresh_mechFlex], and we pattern a large number to build a mechanically reliable anchor. These tethers can be broken individually by micromanipulators while allowing the membranes to withstand harsh processing. Without this design, frames would collapse or delaminate, precluding scalable, high-yield fabrication.
@@
We use X-ray photoelectron spectroscopy to verify surface cleanliness prior to performing a thermal oxygen anneal at 450C according to the procedure described in [Sangtawesin_SurfNoise], which ensures high-quality oxygen termination. We image the resulting NV centers in a confocal microscope (NA = 0.9) under 520 nm illumination. The confocal microscope image overlaid with a scanning electron microscopy (SEM) image of the devices shows bright, single NV center emission in the nanobeams (Figure 1f). Single NV centers are brighter in the nanobeams compared to the host diamond, while their density is similar.
Figure 2: (a) Hahn echo coherence time T2 as a function of NV center depth in nanobeams and the host diamond, plotted together with data from Ref [Sangtawesin_SurfNoise] for shallow NV centers after the oxygen surface termination procedure including a 1200C and only a 800C anneal (as in the current work). Yellow dashed line shows a $T_{2}=kd^{n}+T_{2,0}$ fit to the 800C annealed data from Ref. [Sangtawesin_SurfNoise] ($k=0.047\pm 0.04,n=2.35\pm 0.25,T_{2,0}=3.35\pm 3.87$). Grey shaded area marks the error margin for $n$ with fixed $k=0.047$ and $T_{2,0}=3.35$ showing that most of the data that we measure in the nanobeams falls within this scaling range. (b) T2 measured using an XY-N pulse sequence as a function of the number of pulses, N, (N = 1 is a Hahn echo experiment) for two NV centers in the nanobeams. Dashed line shows $\propto{N^{2/3}}$ dependence. (c) Spectral noise density extracted from a XY-8 decoherence decay using Equation S2 for the same NV centers as in (b). The expected peak position associated with the 13C Larmor frequency is marked with a vertical dashed line. (d) $T_{1,SQ}$ and $T_{1,DQ}$ as a function of NV center depth in nanobeams (purple) and unfabricated diamond (blue).
@@ 2.2 NV Center Spin Properties
We examine the spin properties of NV centers implanted within nanometers of the diamond surface in nanobeams to evaluate their utility as nanoscale quantum sensors, and we compare them to those in the unfabricated regions of the diamond. We measure the Hahn echo coherence time ($T_{2}$) as a function of depth from the surface for several NV centers in the nanobeams and in the unfabricated diamond (Figure 2a). The depth is measured using proton nuclear magnetic resonance, as has been described previously. [pham2016nmr] The $T_{2}$ decreases as NV centers are brought closer to the diamond surface, as has been reported in the literature. [Sangtawesin_SurfNoise, myers2014probing] The observed depth dependence of $T_{2}$ for NV centers in nanobeams is within the distribution of NV centers in the unfabricated regions, and is also consistent with state-of-the-art surface termination as described in the literature. [Sangtawesin_SurfNoise] We note that the coherence times are slightly shorter than the best reported NV centers (Figure 2a, orange points), which resulted from a 1200C anneal that removes multivacancy clusters that form because of implantation damage, pointing to a potential direction for future improvement in the NV center performance in nanobeams.
@@
To probe higher frequency noise, we measure the spin coherence with dynamical decoupling by XY-N pulse sequences. The coherence time $T_{2}$ increases with the number of pulses in the sequence, $N$, for two randomly selected NV centers (Figure 2b). We do not observe any saturation in $T_{2}$ up to 96 pulses, however the scaling deviates from the $N^{2/3}$} scaling expected for a slowly fluctuating noise bath,[bar2012suppression] which can be attributed to the presence of broadband noise induced by defects at the diamond surface. Spectral noise decomposition using an XY-8 pulse sequence[romach2015spectroscopy] for the same two NV centers reveals a noise spectrum $S(\omega)$ consistent with a broadband noise bath and a pronounced peak corresponding to the 13C nuclear spin Larmor precession at 360 G (Figure 2c). These observations are consistent with previous characterization of the surface noise bath in state-of-the-art shallow NV centers. [Sangtawesin_SurfNoise]
@@
Finally, we measure single-quantum spin lifetime ($T_{1,SQ}$) and double-quantum spin lifetime ($T_{1,DQ}$) to probe magnetic and electric noise near the NV center spin resonance frequency ($\approx$ 2.8 GHz)[myers2014probing], correspondingly, as a function of distance from the surface (Figure 2d). The spin lifetimes display a large distribution from 100 µs to several ms, with no clear depth dependence, and the distributions between NV centers in nanobeams and in the unfabricated region are similar. The $T_{1,DQ}$ values between 100 µs and 1 ms are similar to the results presented in [Sangtawesin_SurfNoise], which confirms that the fabrication process does not contribute additional sources of electric noise from the diamond surface. At the same time, we observe several shallow NV centers with $T_{1,SQ}$ in the range of 200 µs to 1 ms compared to $T_{1,SQ}$ ¿ 1 ms as reported in [Sangtawesin_SurfNoise], which could point to slightly increased high-frequency magnetic noise at the surface.